OpenStudy (anonymous):

Brand A cereal costs $6 for 18 ounces. Brand B has an equivalent price. Which of the following could be the cost of Brand B? $5 for 16 ounces $9 for 24 ounces $4 for 10 ounces $8 for 24 ounces

OpenStudy (anonymous):

6/18 = 1/3 8/24 = 1/3

OpenStudy (anonymous):

$8 for 24 ounces

OpenStudy (anonymous):

Yeah that looks about right so I would go with option D as well.

OpenStudy (anonymous):

I got it right, but no one gave me an explanation

OpenStudy (anonymous):

@amistre64 no one gave me an explanation in the question

OpenStudy (amistre64):

then they should all be flogged within inches of their lives :) rob gave an explanantion, it just wasnt all that wordy.

OpenStudy (amistre64):

6 to 18 is a ratio of 1 to 3 thats the same reasoning behind robs reply

OpenStudy (anonymous):

yeah,I mean it should have had more word and should have gave an explanation

OpenStudy (amistre64):

math and words dont mix that well :)

OpenStudy (anonymous):

oohh okay thanks